# Jiwen Yu1∗†  Yiran Qin1∗   Xintao Wang2‡   Pengfei Wan2  Di Zhang2  Xihui Liu1‡

### v1

@@ 3.6 Autoregressive Generation of Long Action-Controllable Game Videos
Cross Attention Cross Attention 8.67 0.3313 107.13 20.46 0.3137 125.67 325.18 0.3103 167.37
Concat Concat 22.37 0.3277 103.89 19.18 0.3159 133.42 258.93 0.3123 145.74
Cross Attention Concat 7.79 0.3292 105.28 18.64 0.3184 127.84 249.54 0.3107 139.91
@@
Current action-controllable video generation methods are limited to fixed-length outputs, which is inadequate for practical game applications requiring continuous video streams. To address this, we develop an autoregressive approach that generates multiple frames per step based on previous outputs, enabling efficient long video generation.
One established approach for autoregressive video generation follows next-token prediction [23, 47, 39], but lacks accessible strong model parameters. While video diffusion t transformer [26, 27, 24, 48] offers superior generation quality, it typically operates in a full-sequence manner. Inspired by Diffusion Forcing [8], which unlike standard diffusion models that require identical noise levels across frames, allows different noise levels where later frames with more noise can depend on earlier frames with less noise, we modify the video diffusion transformer to enable autoregressive generation. As shown in Fig. 6 (a), during training with $N+1$ frame latents (from 00 to $N$), we randomly select the first $k+1$ frames as conditions without adding noise, while adding noise only to the remaining $N-k$ frames for noise prediction training. Although the first $k+1$ frames are input to the model, since they are assumed to be already generated, their predicted noise outputs are not utilized. To improve training efficiency, we focus on computing training losses only for the $N-k$ frames that require prediction. As for inference, as shown in Fig. 6 (b), after full-sequence generation of the first $N+1$ frame latents, we can autoregressively generate new $N-k$ frames by selecting the most recent $k+1$ frames from history video latents as conditions for each subsequent generation, and merge them into the history video latents. This process can be repeated to achieve infinite-length video generation. Unlike conventional next-frame generation methods [8, 36, 11], our approach supports multi-frame generation in one step, greatly reducing the time for long video generation.
Implementation Details.
Since the model structure remains unchanged and only the noise levels vary across frames, the training configuration remains identical to full-sequence generation, including action control. We also add small noise perturbations (equivalent to step $15$/$1000$) to conditional frames, which helps reduce error accumulation in long-term generation [36, 11]. The long video results for action control can be found on our project page.
@@ 4.1 Implementation Details
Figure 7: Demonstration of action control capabilities in the Minecraft domain. The model has successfully learned basic atomic actions (WASD keys) and mouse-based yaw and pitch controls. Additionally, it can combine these atomic actions to execute more complex movements. Note that the text below each video frame is a descriptive label of the content, not a text prompt provided to the model.
Pretrained Model Setting.
Our experiments are based on an internal $1$B-sized transformer-based text-to-video diffusion model for research purpose, which is distilled from a larger pretrained video diffusion model and possesses strong generative priors in the open domain. The resolution of the game videos is $360\times 640$. The temporal compression rate of the VAE is $r=4$.
Training and Inference Setting.
Each phase of fine-tuning or training requires about two-four days of training on $8$ A100 GPUs with a batch size of $64$. The hyper parameters for LoRA finetuning can be referenced as $rank=128$, with a learning rate of 1e-4. The learning rate for training the action control module is 1e-5. We only apply classifier free guidance [18] to the text prompt conditional input and use DDIM [33] sampling with $50$ sampling steps.
Evaluation Setting.
We retained $5\%$ of our collected, segmented dataset as a test set, excluding it from training, and selected three subsets from it for ablation study: (1) only-key: contains only keyboard actions, designed to test the model’s ability to follow discrete actions; (2) mouse-small: includes small-scale continuous mouse movements; and (3) mouse-large: includes large-scale continuous mouse movements. Additionally, for qualitative experiments, we support custom combinations of input actions, allowing us to test complex or rare action combinations. We use the following evaluation metrics: (1) Flow-MSE: calculates the optical flow of the generated video to reflect its dynamics, assessing action-following ability by measuring mean square error to the optical flow of the reference video; (2) CLIP-Sim: computes feature similarity in the CLIP [31] space to evaluate semantic relevance to the given text prompt; (3) FID: measures distribution differences between the generated videos and the reference, providing an assessment of generation quality.
@@ 4.2 Action Controllability
Ablation Study.
We perform ablation studies on the control mechanisms for continuous mouse movement signals and discrete keyboard control signals, comparing two typical methods: cross-attention and concatenation-based injection. The results are presented in Tab. 1. To evaluate how well the model’s generated dynamics align with action controls, we use Flow-MSE as a metric. For discrete control signals, such as keyboard inputs, cross-attention outperforms concatenation, suggesting that category-based signal control benefits from similarity-based cross-attention, as is often used in text-based control. In contrast, for continuous mouse movement signals, concatenation is more effective than cross-attention. This may be due to cross-attention’s similarity computation, which tends to reduce the influence of the control signal’s magnitude, thereby affecting the final result. Additionally, the Flow-MSE values reveal that mouse movements have a larger impact on the visual output than keyboard inputs, particularly in the mouse-large test set where movement magnitude is greater. For the CLIP-Sim and FID metrics, there is little difference between the methods. This is primarily because our multi-phase training strategy decouples style learning into Phase #1, where style learning is consistent across all methods. As a result, metrics assessing semantic consistency and generation quality yield similar performance across different methods.
Demonstration of various action control in Minecraft.
In Fig. 7, we demonstrate the action control capabilities of our proposed action control module within the Minecraft domain. The module successfully learns fundamental atomic actions, such as moving forward, backward, left, and right, as well as adjusting yaw and pitch angles. Additionally, the model learns to combine these actions to achieve more complex control. The diverse scenes in Fig. 7 reflect the scene variety in our dataset, highlighting its strength in capturing a wide range of scenarios. Notably, actions like moving backward, left, and right, although seemingly simple, are rare in typical human Minecraft gameplay. A dataset based on human gameplay would likely suffer from biased distributions, making it challenging to learn these less common actions effectively.
Interaction in generative Minecraft videos.
@@ 4.3 Scene Generalization
In Fig. 1, we showcase a variety of newly generated game videos across open-domain scenes. In theory, any scene within the generative capabilities of the pre-trained model could be included as content in these new games. These results inspire a vision of the future where generative game engines emerge as a new form of gaming, allowing players or game creators to generate and interact with anything they can imagine at minimal cost.
More inspiration from examples of racing games.
In Fig. 9, we discuss an intriguing example. Since our collected Minecraft data is from a first-person perspective, the learned action space primarily generalizes to first-person scenes. Here, we experimented with a racing game scenario using a prompt for a car. Interestingly, we observed that the model’s learned yaw control for the mouse seamlessly generalized to steering control in the racing game. Additionally, certain directional controls, such as moving backward or sideways, were diminished, an adaptation that aligns well with typical controls in a racing game, where these actions are rarely needed. This example not only highlights the strong generalization capabilities of our method but also raises the question: could there exist a larger, more versatile action space that encompasses a wider range of game controls, extending beyond first-person and racing games?
This example also leads us to wonder whether the racing game scenario could have applications in autonomous driving. If we were to collect an action dataset from an autonomous driving simulation environment, could a pre-trained model then generate unlimited open-domain autonomous driving data? Our exploration of scene generalization within generative game engines may hold valuable insights for other fields as well.
5 Potential of Generalizable World Model
@@ 6 Conclusion
Appendix A Details of GF-Minecraft Dataset
@@ A.1 Basic Information
To further enhance the diversity and generalization of the generated game scenes, we preconfigured three biomes (forest, plains, desert), three weather conditions (clear, rain, thunder), and six different times of day (“Starting of a day", “Noon, sun is at its peak", “Sunset", “Beginning of night", “Midnight, moon is at its peak", “Beginning of sunrise"). This process resulted in the generation of 2,000 video clips with action annotations, each containing 2,000 frames.
@@ A.3 Details of Action Space
We use the part of the action space of Minedojo [13] which encompasses nearly all actions available to human players, including keypresses, mouse movements, and clicks. We used keypresses and mouse movements as our control signal. The specific binary actions used in our setup are listed in Table 2. Interface1 to Interface5 represent different MineDojo interfaces, where mutually exclusive actions (e.g., moving forward and backward) are assigned to the same interface. Mouse movements are represented by the offset of the mouse pointer from the center of the game region. For each frame in the video, we calculate the cumulative offset relative to the first frame, and this absolute offset is used as input to the model.
@@
Appendix B Supplementary Comparative Experimental Results
B.1 Qualitative Results of Ablation Study on Action Control Module
Based on the experimental results shown in Table 1, it can be observed that cross attention performs better in terms of action control capability for keyboard inputs, while concatenation performs relatively poorly. In our experiments, we even found that concatenation sometimes struggles to follow discrete keyboard input signals, which is consistent with the significant differences observed in the metrics. Related qualitative results are shown in Fig. 11.
Additionally, the metric results in Table 1 indicate that concatenation has certain advantages over cross attention in modeling continuous mouse movements. This is somewhat similar to some works on camera pose control [37]. However, upon a closer comparison of the qualitative results, we found it challenging to distinguish the differences between them visually. This is reasonable because, as shown in the quantitative results in Table 1, the difference between concatenation and cross attention is relatively small, even for large-scale variations in mouse movements.
B.2 Evaluation on Multi-Phase Training Strategy for Scene Generalization
To evaluate the scene generalization capability of GameFactory in open-domain scenarios, we used the action sequences from the test set along with open-domain prompts and generated open-domain results using a model trained with multi-phase training. Similarly, we used the action sequences and prompts annotated from the test set to generate Minecraft videos with the same action controls using the model trained with multi-phase training. We then compared the Flow-MSE of both results against the reference videos. As shown in Tab. 3, the differences between the two are minor, indicating that the action control capability learned on Minecraft data has been transferred to open-domain videos to a certain extent.
[…]
Multi-Phase Training Minecraft 43.48 - - -
Multi-Phase Training Open-Domain 54.13 0.7565 0.3181 121.18
One-Phase Training Open-Domain 76.02 0.7345 0.3111 167.79
Next, we compared the generalization capability in open-domain scenarios between multi-phase training and one-phase training (i.e., directly training action control on the original parameters) after replacing the text-to-video parameters with the original parameters. The Domain-Sim metric is calculated by extracting CLIP features from the video frames and comparing their similarity in CLIP space to the CLIP features of corresponding video frames generated by the original model with the same prompt. This metric reflects the domain difference between the generated open-domain action-controlled videos and those generated by the original model. According to the results in Tab. 3, multi-phase training demonstrates better action control capability and achieves higher domain similarity with the original model. Additionally, the generated videos exhibit slightly better quality and alignment with the prompts. The qualitative results are shown in Fig. 12, where it can be observed that one-phase training results in degraded action control capability in open-domain scenarios. Moreover, the generated frames display distortion, leading to an overall decline in video quality.
Figure 11: Qualitative comparison of key input control performance. It can be observed that cross-attention significantly outperforms concatenation in handling discrete key input signals, while concatenation may fail to respond to the key input. The yellow buttons indicate pressed keys.

### v2

@@ 4.3 Autoregressive Generation of Long Action-Controllable Game Videos
Current video generation methods introduced in Sec. 3 are limited to fixed-length outputs, which is inadequate for practical game applications requiring continuous video streams. To address this, we develop an autoregressive approach that generates multiple frames per step based on previous outputs, enabling efficient long video generation.
While video diffusion transformer [31, 33, 26, 52] offers superior generation quality, it typically operates in a full-sequence manner. Inspired by Diffusion Forcing [8], we modify the video diffusion transformer to enable autoregressive generation. Specifically, unlike standard diffusion models that require identical noise levels across frames, our approach allows different noise levels where later frames can have more noise while depending on earlier frames with less noise. This varying noise schedule ensures that earlier frames are generated first, allowing subsequent frames to be conditioned on them in an autoregressive manner.
As shown in Fig. 5 (a), during training with $N+1$ frame latents (from 00 to $N$), we randomly select the first $k+1$ frames as conditions without adding noise, while adding noise only to the remaining $N-k$ frames for noise prediction training. Although the first $k+1$ frames are input to the model, since they are assumed to be already generated, their predicted noise outputs are not utilized. To improve training efficiency, we focus on computing training losses only for the $N-k$ frames that require prediction. As for inference, as shown in Fig. 5 (b), after full-sequence generation of the first $N+1$ frame latents, we can autoregressively generate new $N-k$ frames by selecting the most recent $k+1$ frames from history video latents as conditions for each subsequent generation, and merge them into the history video latents. This process can be repeated to achieve infinite-length video generation. Unlike conventional next-frame generation methods [8, 42, 11], our approach supports multi-frame generation in one step, greatly reducing the time for long video generation.
5 Open-Domain Game Scene Generalization
This section introduces how to generalize learned action control capabilities to open-domain scenes. As analyzed in Sec. 1, while pre-trained video models offer rich generative priors for open-domain generation, directly fine-tuning them with game data leads to undesired style bias: the outputs inherit the visual style of training data while learning action control. To address this, we propose a Style-Action Decoupled Model with Domain Adapter (Sec. 5.1) and a multi-phase training strategy (Sec. 5.2).
5.1 Style-Action Decoupling with Domain Adapter
To prevent action control capabilities from being bound to specific game styles, our key insight is to disentangle the learning of game style and action control through different modules and parameters. Specifically, while using the proposed action control module to learn action controllability, we employ an independent domain adapter to capture game-specific visual styles. The domain adapter is implemented using LoRA [21], which efficiently learns specific styles and can be plugged in and out without affecting the open-domain generation priors of the original model. To effectively train these decoupled components, we need a carefully designed training strategy that ensures the style adaptation and action control learning remain independent by learning them in different training phases. We detail this multi-phase training approach in the next section.
5.2 Multi-Phase Training Strategy
DIAMOND [2] GameNGen [42] GameGenX [7] Oasis [11] Matrix [14] Genie 2 [12] GameFactory
Release Time NeurIPS 2024 ICLR 2025 ICLR 2025 2024.10.31 2024.12.4 2024.12.4 -
Game Sources Atari, CS:GO DOOM AAA Games Minecraft AAA Games Unknown Minecraft
Resolution $280\times 150$ $240$p $720$p $640\times 360$ $720$p $720$p $640\times 360$
Control Granularity Frame-level Frame-level Video-level Frame-level Frame-level Frame-level Frame-level
Technical Paper
Testable Model -
Available Dataset
Action Space 18 Keys Key Instruction Key + Mouse 4 Keys Key+Mouse 7 Keys+Mouse
Scene Generalizable
Table 1: Comparison with recent related works. Scene-generalizable action control is the core contribution of GameFactory, being the only technical paper that validates on complex action space and conducts extensive testing across open-domain scenarios.
❑ Phase #1: Tune LoRA to Fit Game Videos. We fine-tune the pre-trained video diffusion model using LoRA to adapt it to specific game video while preserving most original parameters. This produces a model specialized for the target game domain. Better style adaptation in this phase allows the next phase to focus purely on action control, reducing style-control entanglement.
❑ Phase #2: Tune Action Control Module. We freeze both pre-trained parameters and LoRA, only training the action control module with game videos and action signals. Since Phase #1 has handled style adaptation through LoRA, the training loss now focuses less on style learning. This allows the model to concentrate on action control learning, as it becomes the main contributor to minimizing the diffusion loss. Such separation enables style-independent control that can generalize to open-domain scenarios.
❑ Phase #3: Inference on Open Domain. During inference, we remove the LoRA weights for game style adaptation, keeping only the action control module parameters. Thanks to the decoupling in previous phases, the action control module can now work independently of specific game styles, enabling controlled game video generation across open-domain scenarios.
@@ 5.2 Multi-Phase Training Strategy
Cross-Attn Cross-Attn 0.0527 8.67 0.3313 107.13 814.05 0.0798 20.46 0.3137 125.67 1203.29 0.1362 325.18 0.3103 167.37 1383.92
Concat Concat 0.0853 22.37 0.3277 103.89 786.50 0.0756 19.18 0.3159 133.42 1151.71 0.1179 258.93 0.3123 145.74 1405.47
Cross-Attn Concat 0.0439 7.79 0.3292 105.28 795.03 0.0685 18.64 0.3184 127.84 1032.98 0.1021 249.54 0.3107 139.91 1420.89
@@ 6.1 Implementation Details
❑ Pretrained Model Setting. Our experiments are based on an internal $1$B-sized transformer-based text-to-video diffusion model for research purpose, which is distilled from a larger pretrained video diffusion model and possesses strong generative priors in the open domain. The resolution of the game videos is $360\times 640$. The temporal compression rate of the VAE is $r=4$.
❑ Training and Inference Setting. Each phase of fine-tuning or training requires about two-four days of training on $8$ A100 GPUs with a batch size of $64$. The hyper parameters for LoRA finetuning can be referenced as $rank=128$, with a learning rate of 1e-4. The learning rate for training the action control module is 1e-5. We only apply classifier free guidance [18] to the text prompt conditional input and use DDIM [39] sampling with $50$ sampling steps.
❑ Evaluation Setting. We retained $5\%$ of our collected, segmented dataset as a testset, excluding it from training, and selected three subsets for ablation study: (1) only-key: contains only keyboard actions, designed to test the model’s ability to follow discrete actions; (2) mouse-small: includes small-scale continuous mouse movements; and (3) mouse-large: includes large-scale continuous mouse movements. Additionally, for qualitative experiments, we support custom combinations of input actions, allowing us to test complex or rare action combinations. We use these evaluation metrics: (1) Flow: calculates the optical flow of the generated video to reflect its dynamics, assessing action-following ability by measuring mean square error to the optical flow of the reference video; (2) Cam: computes the Euclidean distance between camera poses extracted from predicted videos and those extracted from reference videos, where both sets of camera poses are obtained using GLOMAP [32]. (3) CLIP: computes feature similarity in the CLIP [37] space to evaluate semantic relevance to the given text prompt; (4) FID, FVD: measures distribution differences between the generated videos and the reference, providing an assessment of generation quality.
❑ Comparison with Related Works. As shown in Tab. 1, given the diverse sources of game datasets, varying video resolutions, and different levels of control across methods, establishing a unified benchmark for comparison is challenging. Regarding our main contribution of scene-generalizable action control, the only capable technical paper [14] validated their approach on a simple action space (left turn, right turn, acceleration in racing games). In contrast, we validate our method on a much more comprehensive action space (forward/backward/left/right movement, jumping, acceleration/deceleration, mouse movement) and provide extensive results in the supplementary materials.
@@ 6.2 Action Controllability
❑ Ablation Study.
Figure 7: Qualitative comparison of key input control performance. The yellow buttons indicate pressed keys.
[…]
Multi-Phase In- 0.0839 43.48 - - - -
Multi-Phase Open- 0.0997 54.13 0.7565 0.3181 121.18 1256.94
One-Phase Open- 0.1134 76.02 0.7345 0.3111 167.79 1323.58
We perform ablation studies on the control mechanisms for continuous mouse movement and discrete keyboard control, comparing two typical methods: cross-attention and concatenation. The results are presented in Tab. 2 and Fig. 7. For discrete keyboard inputs, cross-attention outperforms concatenation. As shown in Fig. 7, concatenating the action control signal with the input leads to poor action following performance. This suggests that category-based signal control benefits from similarity-based cross-attention, as is often used in text-based control. In contrast, for continuous mouse movement, concatenation is more effective than cross-attention. This may be due to cross-attention’s similarity computation, which tends to reduce the influence of the control signal’s magnitude, thereby affecting the final result. Additionally, the values of Cam and Flow reveal that mouse movements have a larger impact on the visual output than keyboard inputs, particularly in the mouse-large test set where movement magnitude is greater. For the metrics of CLIP, FID and FVD, there is little difference between the methods. This is primarily because our multi-phase training strategy decouples visual style learning into Phase #1, where style learning is consistent across all methods. As a result, metrics assessing semantic consistency and generation quality yield similar performance across different methods.
@@ 6.3 Scene Generalization
❑ Create new games in generalized scenes. In Fig. 1, we showcase a variety of newly generated game videos across open-domain scenes. These results inspire a future where generative game engines emerge as a new form of gaming, allowing players or game creators to generate and interact with anything they can imagine at minimal cost.
❑ Comparison Results. As shown in Tab. 3, we evaluate the action following capability using Cam and Flow metrics on in-domain as baseline. The results demonstrate that our multi-phase training strategy achieves better action following performance closer to the baseline compared to the one-phase approach. Furthermore, the multi-phase strategy shows superior performance across text-alignment and generation quality metrics including CLIP, FID, and FVD. The Dom metric, which measures the CLIP space similarity between videos generated by the original model and the fine-tuned model, indicates that the multi-phase trained model maintains a domain closer to the original model. This is visually confirmed in Fig. 8, where the multi-phase strategy preserves the original model’s generation domain without style leakage. Additionally, the one-phase approach degrades the generated video quality with noticeable artifacts.
6.4 Evaluation for GF-Minecraft Dataset
[…]
VPT [3] 0.1324 107.67 0.3174 156.69 1233.15
GF-Minecraft (ours) 0.0839 43.48 0.3135 125.85 1047.59
Table 4: Comparison with Minecraft dataset with human bias.
Dataset W A S D Space Shift Ctrl
VPT [3] 50.11% 4.03% 0.32% 3.45% 20.37% 0.14% 19.58%
Ours 13.56% 13.56% 13.56% 13.56% 15.25% 15.25% 15.25%
Table 5: The proportion of key inputs across different datasets.
Figure 9: Compare the dataset on actions that are less commonly used by human players to test the effect of human bias in dataset.
In Sec. 4.1, we introduced our GF-Minecraft dataset, which eliminates human bias by ensuring uniform distribution of atomic actions. This design enables models trained on GF-Minecraft to effectively respond to actions that human players rarely perform. We compare our dataset with the VPT [3] dataset, which contains recordings of human gameplay and inherent human biases. Specifically, we selected video clips from VPT’s Find Cave dataset, as it closely matches our setting by primarily excluding inventory management and block modification operations. Tab. 5 compares the usage proportions of keyboard inputs across both datasets. The VPT dataset shows highly skewed distributions. For instance, the forward movement key (W) appears over 100 times more frequently than the backward movement key (S), reflecting typical human gameplay patterns. Tab. 4 compares the action control capabilities of models trained on both datasets through in-domain evaluation, demonstrating clear advantages of our GF-Minecraft dataset in action following performance. Fig. 9 illustrates two representative examples: jumping in place and moving backward (both actions rarely performed by human players). The model trained on GF-Minecraft successfully follows these actions, while the VPT-trained model fails to execute them properly. Specifically, when commanded to jump in place, the VPT model incorrectly combines forward movement with jumping, and when instructed to move backward, it simply remains stationary.
6.5 Evaluation for Long Video Generation
[…]
All frames 0.1547 148.73 0.2965 176.07 1592.43
Only predicted frames 0.0924 85.45 0.3190 136.95 1154.45
Table 6: Ablation study on loss scope for long video generation.
Figure 10: Demonstration of key frames in generated long video.
In Tab. 6, we compare different scopes for computing loss functions in long video generation training: calculating losses across all frames versus only on frames that need to be predicted. The results demonstrate that computing losses solely on frames requiring prediction yields better performance. This improvement can be attributed to the elimination of noise interference from previously generated frames, as learning from such noise is irrelevant for future video generation. As demonstrated in Fig. 5, our model successfully generates long videos exceeding 300 frames in length. Additional examples of long video generation can be found in the supplementary materials.
@@ 7 Conclusion
Appendix A Details of Minecraft Dataset
@@ A.1 Basic Information
To further enhance the diversity and generalization of the generated game scenes, we preconfigured three biomes (forest, plains, desert), three weather conditions (clear, rain, thunder), and six different times of day ("Starting of a day," "Noon, sun is at its peak," "Sunset," "Beginning of night," "Midnight, moon is at its peak," "Beginning of sunrise"). This process resulted in the generation of 2,000 video clips with action annotations, each containing 2,000 frames.
@@ A.3 Details of Action Space
We use the part of the action space of Minedojo [13] which encompasses nearly all actions available to human players, including keypresses, mouse movements, and clicks. We used keypresses and mouse movements as our control signal. The specific binary actions used in our setup are listed in Tab 7. Interface1 to Interface5 represent different MineDojo interfaces, where mutually exclusive actions (e.g., moving forward and backward) are assigned to the same interface. Mouse movements are represented by the offset of the mouse pointer from the center of the game region. For each frame in the video, we calculate the cumulative offset relative to the first frame, and this absolute offset is used as input to the model.
@@
Appendix B Supplementary Experimental Results
B.1 Interaction in generative Minecraft videos.
B.2 More Inspiration from Examples of Racing Games
In Fig. 13, we discuss an intriguing example. Since our collected Minecraft data is from a first-person perspective, the learned action space primarily generalizes to first-person scenes. Here, we experimented with a racing game scenario using a prompt for a car. Interestingly, we observed that the model’s learned yaw control for the mouse seamlessly generalized to steering control in the racing game. Additionally, certain directional controls, such as moving backward or sideways, were diminished, an adaptation that aligns well with typical controls in a racing game, where these actions are rarely needed. This example not only highlights the strong generalization capabilities of our method but also raises the question: could there exist a larger, more versatile action space that encompasses a wider range of game controls, extending beyond first-person and racing games? This example also leads us to wonder whether the racing game scenario could have applications in autonomous driving. If we were to collect an action dataset from an autonomous driving simulation environment, could a pre-trained model then generate unlimited open-domain autonomous driving data? Our exploration of scene generalization within generative game engines may hold valuable insights for other fields as well.
Appendix C Potential of Generalizable World Model